A firm operates in a perfectly competitive market. The market price of its product is 4
1
birr and the total cost function is given by TC= Q 3  5Q 2  20Q  50 , where TC is the
3
total cost and Q is the level of output.
A) What level of output
should the firm produce to
maximize its profit?
B) Determine the level of profit at
equilibrium.
C) Find TR & TC
D) What minimum price is
required by the firm to stay
in the market?
1
Expert's answer
2020-08-03T15:11:30-0400

P = 4 birr, "TC = 1\/3*Q^3 - 5Q^2 + 20Q + 50."

A) The firm produce at P = MC to maximize its profit:

"MC = Q^2 - 10Q + 20,"

"Q^2 - 10Q + 20 = 4,"

"Q^2 - 10Q + 16 = 0,"

Q1 = 8 units, Q2 = 2 units.

B) The level of profit at equilibrium is:

"TP = TR - TC = 4*8 - (1\/3*8^3 - 5*8^2 + 20*8 + 50) = -28.67."

The profit at Q = 2 is lower, so it is not a profit-maximizing quantity.

C) TR = 4*8 = 32,

TC = "1\/3*8^3 - 5*8^2 + 20*8 + 50" = 60.67.

D) The minimum price is required by the firm to stay in the market is:

P = AVC = VC/Q = 1/3*8^2 - 5*8 + 20 = 1.33.